### v3

# Graph Diffusion Transformers for Multi-Conditional Molecular Generation

Abstract
Inverse molecular design with diffusion models holds great potential for advancements in material and drug discovery. Despite success in unconditional molecular generation, integrating multiple properties such as synthetic score and gas permeability as condition constraints into diffusion models remains unexplored. We present the Graph Diffusion Transformer (Graph DiT) for multi-conditional molecular generation. Graph DiT integrates an encoder to learn numerical and categorical property representations with the Transformer-based denoiser. Unlike previous graph diffusion models that add noise separately on the atoms and bonds in the forward diffusion process, Graph DiT is trained with a novel graph-dependent noise model for accurate estimation of graph-related noise in molecules. We extensively validate Graph DiT for multi-conditional polymer and small molecule generation. Results demonstrate the superiority of Graph DiT across nine metrics from distribution learning to condition control for molecular properties. A polymer inverse design task for gas separation with feedback from domain experts further demonstrates its practical utility.
1 Introduction
Diffusion models for molecular graphs are essential for inverse design of materials and drugs by generating molecules and polymers (macro-molecules) [thornton2012polymer, wu2018moleculenet], because the models can be effectively trained to predict discrete graph structures and atom/bond types in denoising processes [vignac2022digress]. Practical inverse designs consider multiple factors such as molecular synthetic score and various properties [gebauer2022inverse], known as the task of multi-conditional graph generation.
Existing work converted multiple conditions into a single one and solved the task as single-condition generation [bilodeau2022generative, lee2023exploring]. However, multi-property relations may not be properly or explicitly defined [bilodeau2022generative]. First, the properties have diverse scales and units. For example, the synthetic complexity ranges from 1 to 5 [coley2018scscore], while the gas permeability varies widely, exceeding 10,000 in Barrier units [barnett2020designing]. This gap makes it hard for models to balance the conditions. Second, multi-conditions consist of a mix of categorical and numerical properties. The common practice of addition [xie2021mars] or multiplication [lee2023exploring] is inadequate for combination.
Figure 1(a) empirically illustrates the challenges in multi-conditional generation, i.e., discovering molecules meeting multiple properties. We used a test set of 100 data points with three properties: synthesizability (Synth.) [ertl2009estimation], O2 and N2 permeability (O2Perm and N2Perm) [barnett2020designing]. A single-conditional diffusion model generated up to 30 graphs for each condition, resulting in a total of 90 graphs for three conditions. We sort the 30 graphs in each set using a polymer property Oracle (see LABEL:add:subsec:motivation-setup). Then, we check whether a shared polymer structure that meets multi-property constraints can be identified across different condition sets. If we find the polymer, its rank $K$ (where $K$ is between 1 and 30) indicates how high it appears on the lists, considering all condition sets. If not, we set $K$ as 30. Figure 1(a) shows the frequency distribution of $K$ on the 100 test cases. The median $K$ was 30, indicating that the multiple properties were not met on over half of the test polymers despite generating a large number of graphs.
(a) Existing work’s limitation: A median rank of 30 showed that on fewer than half test polymers, the sets of generated graphs from different single conditions intersected, indicating a failure to generate polymers meeting multiple properties.
(b) Proposed work: Our idea, multi-conditional guidance for diffusion models, successfully generated polymers that satisfied multi-property constraints. It achieved a higher rank than 30 in any set of the single-conditional generated graphs.
Figure 1: Multi-conditional diffusion guidance in (b) generates polymers of higher property accuracy than existing work in (a). Explanations are in Section 1 and details are in LABEL:add:subsec:motivation-setup.
To address these challenges, we project multi-properties into representations by learning, thereby guiding the diffusion process for molecule generation. We propose the Graph Diffusion Transformer (Graph DiT) for graph denoising under conditions. Graph DiT has a condition encoder for property representation learning and a graph denoiser. The condition encoder utilizes a novel clustering-based method for numerical properties and one-hot encoding for categorical ones to learn multi-property representations. The graph denoiser first integrates node and edge features into graph tokens, then denoise these tokens with adaptive layer normalization (AdaLN) in Transformer layers [huang2017arbitrary, peebles2023scalable]. AdaLN replaces the molecular statistics (mean and variance) in each hidden layer with those from the condition representation, effectively outperforming other predictor-based and predictor-free conditioning methods [jo2022score, vignac2022digress, peebles2023scalable], as shown in LABEL:subsec:ablation. We observe that existing forward diffusion processes [vignac2022digress, jo2022score] apply noise separately to atoms and bonds, which may compromise the accuracy of Graph DiT in noise estimation. Hence, we propose a novel graph-dependent noise model that effectively applies noise tailored to the dependencies between atoms and bonds within the graph.
Results in Figure 1(b) show that the polymers generated by Graph DiT closely align with multi-property constraints. For each test case, we have one graph generated from Graph DiT conditional on three properties. The Oracle determines the rank of this graph among 30 single-conditionally generated graphs for each condition. We find the median ranks are 4, 9, and 11, for Synth., O2 Perm, and N2 Perm, respectively, all much higher than 30. Note that the ranked set of 30 graphs was very competitive because the model was trained on the specific condition dedicatedly.
In experiments, we evaluate model performance on one polymer and three small molecule datasets. The polymer dataset includes four numerical conditions for multi-conditional evaluation. Our model has the lowest average mean absolute error (MAE), significantly reducing the error by 17.86% compared to the best baseline. It also excels in small molecule tasks, achieving over 0.9 accuracy on task-related categorical conditions, notably surpassing the baseline accuracy of less than 0.6. We also examine the model’s utility in inverse polymer designs for O2/N2 gas separation, with domain expert feedback highlighting our model’s practical utility in multi-conditional molecular design.
2 Problem Definition
2.1 Multi-Conditional Inverse Molecular Design
A molecular graph $G=(V,E)$ consists of a set of nodes (atoms) $V$ and edges (bonds) $E$. We follow [vignac2022digress] and define “non-bond” as a type of edge. There are $N$ atoms and each atom has a one-hot encoding, denoting the atom type. We represent it as $\mathbf{X}_{V}\in\mathbb{R}^{N\times F_{V}}$, where $F_{V}$ is the total number of atom types. Similarly, the bond features are a tensor $\mathbf{X}_{E}\in\mathbb{R}^{N\times N\times F_{E}}$, representing both the graph structure and $F_{E}$ bond types.
Let $\mathcal{C}=\{c_{1},c_{2},\dots,c_{M}\}$ be a set of $M$ numerical and categorical conditions. The task is: $q(G\mid c_{1},c_{2},\dots,c_{M})\propto q(G)q(c_{1},c_{2},\dots,c_{M}\mid G)$, where $q$ represents observed probability. We use a model parameterized by $\theta$ for multi-conditional molecular generation $p_{\theta}(G\mid\mathcal{C})$. The evaluation involves both distribution learning $q(G)$ [polykovskiy2020molecular] and condition control $q(c_{1},c_{2},\dots,c_{M}\mid G)$. We follow previous work in assuming that there exist different oracle functions $\mathcal{O}$ that can independently evaluate each conditioned property [gao2022sample]: $q(c_{1},c_{2},\dots,c_{M}\mid G)=\prod_{i=1}^{M}\mathcal{O}_{i}(c_{i}\mid G)$. Note that the oracles are not used in the training of $p_{\theta}$.
2.2 Diffusion Model on Graph Data
Diffusion models consist of forward and reverse diffusion processes [ho2020denoising]. We refer to the forward diffusion process as the diffusion process following [ho2020denoising]. The diffusion process $q(G^{1:T}\mid G^{0})=\prod_{t=1}^{T}q(G^{t}\mid G^{t-1})$ corrupts molecular graph data ($G^{0}=G$) into noisy states $G^{t}$. As timesteps $T\rightarrow\infty$, $q(G^{T})$ converges a stationary distribution $\pi(G)$. The reverse Markov process $p_{\theta}(G^{0:T})=q(G^{T})\prod_{t=1}^{T}p_{\theta}(G^{t-1}\mid G^{t})$, parameterized by neural networks, gradually denoises the latent states toward the desired data distribution.
Diffusion Process
One may perturb $G^{t}$ in a discrete state-space to capture the structural properties of molecules [vignac2022digress]. Two transition matrices $\mathbf{Q}_{V}\in\mathbb{R}^{F_{V}\times F_{V}}$ and $\mathbf{Q}_{E}\in\mathbb{R}^{F_{E}\times F_{E}}$ are defined for nodes $\mathbf{X}_{V}$ and edges $\mathbf{X}_{E}$, respectively [vignac2022digress]. Then, each step $q(G^{t}\mid G^{t-1},G^{0})=q(G^{t}\mid G^{t-1})$ in the diffusion process is sampled as follows.
[…]
where $\operatorname{Cat}(\mathbf{X};\mathbf{p})$ denotes sampling from a categorical distribution with probability $\mathbf{p}$. We remove the subscript (V/E) when the description applies to both nodes and edges. It is assumed that the noise $\mathbf{Q}^{i}$ ($i\leq t$) is independently applied to $\mathbf{X}$ in each step $i$, allowing us to rewrite $q(\mathbf{X}^{t}\mid\mathbf{X}^{t-1})$ as the probability of the initial state $q(\mathbf{X}^{t}\mid\mathbf{X}^{0})=\operatorname{Cat}\left(\mathbf{X}^{t};\mathbf{p}=\mathbf{X}^{0}\bar{\mathbf{Q}}^{t}\right)$, where $\bar{\mathbf{Q}}^{t}=\prod_{i\leq t}\mathbf{Q}^{i}$.
Noise Scheduling
Transition matrices $\mathbf{Q}_{V}$ and $\mathbf{Q}_{E}$ control the noise applied to atom features and bond features, respectively. vignac2022digress defined $\pi(G)=(\mathbf{m}_{X}\in\mathbb{R}^{F_{V}},\mathbf{m}_{E}\in\mathbb{R}^{F_{E}})$ as the marginal distributions of atom types and bond types. The transition matrix at timestep $t$ is $\mathbf{Q}^{t}=\alpha^{t}\mathbf{I}+(1-\alpha^{t})\mathbf{1}\mathbf{m}^{\prime}$ for atoms or bonds, where $\mathbf{m}^{\prime}$ denotes the transposed row vector. Therefore, we have $\bar{\mathbf{Q}}^{t}=\bar{\alpha}^{t}\mathbf{I}+(1-\bar{\alpha}^{t})\mathbf{1}\mathbf{m}^{\prime}$, where $\bar{\alpha}^{t}=\prod_{\tau=1}^{t}\alpha^{\tau}$. The cosine schedule [nichol2021improved] is often chosen for $\bar{\alpha}^{t}=\cos(0.5\pi(t/T+s)/(1+s))^{2}$.
Reverse Process
With the initial sampling $G^{T}\sim\pi(G)$, the reverse process generates $G^{0}$ iteratively in reversed steps $t=T,T-1,\dots,0$. We use a neural network to predict the probability $p_{\theta}(\tilde{G}^{0}\mid G^{t})$ as the product over nodes and edges [austin2021structured, vignac2022digress]:
$p_{\theta}(\tilde{G}^{0}\mid G^{t})=\prod_{v\in V}p_{\theta}(v^{t-1}\mid G^{t})\prod_{e\in E}p_{\theta}(e^{t-1}\mid G^{t})$ (2)
$p_{\theta}(\tilde{G}^{0}\mid G^{t})$ could be combined with $q(G^{t-1}\mid G^{t},G^{0})$ to estimate the reverse distribution on the graph $p_{\theta}(G^{t-1}\mid G^{t})$. For example, $p_{\theta}(v^{t-1}\mid G^{t})$ is marginalized over predictions of node types $\tilde{v}\in\tilde{\mathbf{x}}_{v}$, which applies similarly to edges:
$p_{\theta}(v^{t-1}\mid G^{t})=\sum_{\tilde{v}\in\tilde{\mathbf{x}}_{v}}q(v^{t-1}\mid\tilde{v},G^{t})p_{\theta}(\tilde{v}\mid G^{t}).$ (3)
The neural network could be trained to minimize the negative log-likelihood [vignac2022digress].
$L=\mathbb{E}_{q(G^{0})}\mathbb{E}_{q(G^{t}|G^{0})}\left[-\mathbb{E}_{\mathbf{x}\in G^{0}}\log p_{\theta}\left(\mathbf{x}\mid G^{t}\right)\right]$ (4)
where $\mathbf{x}\in G^{0}$ denotes the node or edge features. Typically, the reverse process in diffusion models does not consider molecular properties as conditions. While there have been efforts to introduce property-related guidance using additional predictors, the more promising approach of predictor-free guidance [ho2022classifier], particularly in multi-conditional generation, remains underexplored.
3 Multi-Conditional Graph Diffusion Transformers
We present the denoising framework of Graph DiT in Figure 2. The condition encoder learns the representation of $M$ conditions. The statistics of this representation like mean and variance are used to replace the ones from the molecular representations [huang2017arbitrary] (see Section 3.2). Besides, we introduce a new noise model in the diffusion process to better fit graph-structured molecules (see Section 3.1).
Figure 2: Denoising framework and architectures for Graph DiT. Details are in Section 3.2.
3.1 Graph-Dependent Noise Models
The transition probability of a node or an edge should rely on the joint distribution of nodes and edges in the prior state. However, as an example shown in Eq. 1, current diffusion models [jo2022score, vignac2022digress, lee2023exploring] treat node and edge state transitions as independent, misaligning with the denoising process in Eq. 3. This difference between the sampling distributions of noise in the diffusion and reverse processes introduces unnecessary challenges to multi-conditional molecular generations.
[…]
$\mathbf{Q}_{G}=\begin{bmatrix}\mathbf{Q}_{V}&\mathbf{1}_{N}^{\prime}\otimes\mathbf{Q}_{VE}\\ \mathbf{1}_{N}\otimes\mathbf{Q}_{EV}&\mathbf{1}_{N\times N}\otimes\mathbf{Q}_{E}\end{bmatrix},$ (5)
where $\otimes$ denotes the Kronecker product, $\mathbf{1}_{N}$, $\mathbf{1}_{N}^{\prime}$, and $\mathbf{1}_{N\times N}$ represent the column vector, row vector, and matrix with all 1 elements, respectively. According to Eq. 5, the first $F_{V}$ columns in $\mathbf{Q}_{G}$ determine the node feature transitions based on both node features (first $F_{V}$ rows) and edge features (remaining $N\cdot F_{E}$ rows). Conversely, the remaining $N\cdot F_{E}$ columns determine the edge feature transitions, depending on the entire graph. We introduce a new diffusion noise model:
$q(\mathbf{X}_{G}^{t}\mid\mathbf{X}_{G}^{t-1})=\widetilde{\operatorname{Cat}}\left(\mathbf{X}_{G}^{t};\tilde{\mathbf{p}}=\mathbf{X}^{t-1}_{G}\mathbf{Q}_{G}^{t}\right),$ (6)
where $\tilde{\mathbf{p}}$ is the unnormalized probability and $\widetilde{\operatorname{Cat}}$ denotes categorical sampling: The first $F_{V}$ columns of $\tilde{\mathbf{p}}$ are normalized to sample $\mathbf{X}_{V}^{t}$, while the remaining $N\cdot E$ dimensions are reshaped and normalized to sample edges $\mathbf{X}_{E}^{t}$. These components are combined to form $\mathbf{X}_{G}^{t}$, completing the $\widetilde{\operatorname{Cat}}$ sampling.
Choice of $\mathbf{Q}_{VE}$ and $\mathbf{Q}_{EV}$
Similar to the definitions of $\mathbf{m}_{V}$ and $\mathbf{m}_{E}$ [vignac2022digress], we leverage the prior knowledge within the training data for the formulation of task-specific matrices, $\mathbf{Q}_{EV}$ and $\mathbf{Q}_{VE}$. We calculate co-occurrence frequencies of atom and bond types in training molecular graphs to obtain the marginal atom-bond co-occurrence probability distribution. For each bond type, each row in $\mathbf{m}_{EV}$ represents the probability of co-occurring atom types. $\mathbf{m}_{VE}$ is the transpose of $\mathbf{m}_{EV}$ and has a similar meaning. Subsequently, we define $\mathbf{Q}_{EV}=\bar{\alpha}^{t}\mathbf{I}+(1-\bar{\alpha}^{t})\mathbf{1}\mathbf{m}_{EV}^{\prime}$ and $\mathbf{Q}_{VE}=\bar{\alpha}^{t}\mathbf{I}+(1-\bar{\alpha}^{t})\mathbf{1}\mathbf{m}_{VE}^{\prime}$.
3.2 Denoising Models with Multi-Property Conditions
We present Graph DiT as the denoising model to generate molecules under multi-conditions $\mathcal{C}=\{c_{1},c_{2},\dots,c_{M}\}$ without extra predictors.
Predictor-Free Guidance
The predictor-free reverse process $\hat{p}_{\theta}(G^{t-1}\mid G^{t},\mathcal{C})$ aims to generate molecules with a high probability $q(\mathcal{C}\mid{G}^{0})$. This could be achieved by a linear combination of the log probability for unconditional and conditional denoising [ho2022classifier]:
$\begin{split}\hat{p}_{\theta}(G^{t-1}\mid G^{t},\mathcal{C})=\log p_{\theta}(G^{t-1}\mid G^{t})+s\left(\log p_{\theta}(G^{t-1}\mid G^{t},\mathcal{C})-\log p_{\theta}(G^{t-1}\mid G^{t})\right),\end{split}$ (7)
where $s$ denotes the scale of conditional guidance. Unlike classifier-free guidance [ho2022classifier], which typically predicts noise, we directly estimate $p_{\theta}(\tilde{G}^{0}\mid G^{t},\mathcal{C})$. We one one denoising model $f_{\theta}(G^{t},\mathcal{C})$ for both $p_{\theta}(\tilde{G}^{0}\mid G^{t})$ and $p_{\theta}(\tilde{G}^{0}\mid G^{t},\mathcal{C})$. Here, $f_{\theta}(G^{t},\mathcal{C}=\emptyset)$ computes the unconditional probability by substituting the original conditional embeddings with the null value. During training, we randomly drop the condition with a ratio, i.e., $\mathcal{C}=\emptyset$, to learn the embedding of the null value. $f_{\theta}(G^{t}=\mathbf{X}_{G}^{t},\mathcal{C})$ comprises two components: the condition encoder and the graph denoiser. An overview of the architecture is presented in Figure 2.
Condition Encoder
We treat the timestep $t$ as a special condition and follow [nichol2021glide] to obtain a $D$-dimensional representation $\mathbf{t}$ with sinusoidal encoding. For property-related numerical or categorical condition $c_{i}\in\mathcal{C}$, we apply distinct encoding operations to get $D$-dimensional representation. For a categorical condition, we use the one-hot encoding. For a numerical variable, we introduce a clustering encoding method. This defines learnable centroids, assigning $c_{i}$ to clusters, and transforming the soft assignment vector of condition values into the representation. It could be implemented using two $\operatorname{Linear}$ layers and a $\operatorname{Softmax}$ layer in the middle as: $\operatorname{Linear}\left(\operatorname{Softmax}\left(\operatorname{Linear}(c_{i})\right)\right)$. Finally, we could obtain the representation of the condition as $\mathbf{c}=\sum_{i=1}^{M}\operatorname{encode}(c_{i})$, where $\operatorname{encode}$ is the specific encoding method based on the condition type. For numerical conditions, we evaluate our proposed clustering-based approach against alternatives like direct or interval-based encodings [liu2023semi]. As noted in LABEL:subsec:ablation, the clustering encoding outperforms the other methods.
Graph Denoiser: Transformer Layers
Given the noisy graph at timestep $t$, the graph tokens are first encoded into the hidden space as $\mathbf{H}=\operatorname{Linear}(\mathbf{X}_{G}^{t})$, where $\mathbf{H}\in\mathbb{R}^{N\times D}$. We then adapt the standard Transformer layers [vaswani2017attention] with self-attention and multi-layer perceptrons (MLP), but replace the normalization with the adaptive layer normalization ($\operatorname{AdaLN}$) controlled by the representations of the conditions [huang2017arbitrary, peebles2023scalable]: $\mathbf{H}=\operatorname{AdaLN}(\mathbf{H},\mathbf{c}$). For each row $\mathbf{h}$ in $\mathbf{H}$:
$\operatorname{AdaLN}\left(\mathbf{h},\mathbf{c}\right)=\gamma_{\theta}(\mathbf{c})\odot\frac{\mathbf{h}-\mu\left(\mathbf{h}\right)}{\sigma\left(\mathbf{h}\right)}+\beta_{\theta}(\mathbf{c}),$ (8)
where $\mu(\cdot)$ and $\sigma(\cdot)$ are mean and variance values. $\odot$ indicates element-wise product. $\gamma_{\theta}(\cdot)$ and $\beta_{\theta}(\cdot)$ are neural network modules in $f_{\theta}(\cdot)$, each of which consists of two linear layers with $\operatorname{SiLU}$ activation [elfwing2018sigmoid] in the middle. We have a gated variant $\operatorname{AdaLN}_{gate}$ for residuals:
$\operatorname{AdaLN}_{gate}\left(\mathbf{h},\mathbf{c}\right)=\alpha_{\theta}(\mathbf{c})\odot\operatorname{AdaLN}\left(\mathbf{h},\mathbf{c}\right)$ (9)
We apply the zero initialization for the first layer of $\gamma_{\theta}(\cdot),\beta_{\theta}(\cdot)$, and $\alpha_{\theta}(\cdot)$ [peebles2023scalable]. There are other options to learn the structure representation from the condition [peebles2023scalable]: $\operatorname{In-Context}$ conditioning adds condition representation to the structure representation at the beginning of the structure encoder, and $\operatorname{Cross-Attention}$ calculates cross-attention between the condition and structure representation. We observe in LABEL:subsec:ablation that $\operatorname{AdaLN}$ performs best among them.
Graph Denoiser: Final MLP
We have the hidden states $\mathbf{H}$ after the final Transformer layers, the MLP is used to predict node probabilities $\tilde{\mathbf{X}}_{V}^{0}$ and edge probabilities $\tilde{\mathbf{X}}_{E}^{0}$ at $t=0$:
$\tilde{\mathbf{X}}_{G}^{0}=\operatorname{AdaLN}(\operatorname{MLP}(\mathbf{H}),\mathbf{c}).$ (10)
We split the output $\mathbf{X}_{G}$ into atom and bond features $\tilde{\mathbf{X}}_{V}^{0},\tilde{\mathbf{X}}_{E}^{0}$. The first $F_{V}$ dimensions of $\tilde{\mathbf{X}}_{G}^{0}$ represent node type probabilities, and the remaining $N\cdot F_{E}$ dimensions cover probabilities for $N$ edge types associated with the node, as detailed in Section 3.1.
Generation to Molecule Conversion
A common way of converting generated graphs to molecules selects only the largest connected component [vignac2022digress], denoted as Graph DiT-LCC in our model. For Graph DiT, we connect all components by randomly selecting atoms. It minimally alters the generated structure to more accurately reflect model performance than Graph DiT-LCC.
4 Experiment
RQ1: We validate the generative power of Graph DiT compared to baselines from molecular optimization and diffusion models in LABEL:subsec:generation-main. RQ2: We study a polymer inverse design for gas separation in LABEL:subsec:study. RQ3: We conduct further analysis to examine Graph DiT in LABEL:subsec:ablation.
4.1 Experimental Setup
We use datasets with over ten types of atoms and up to fifty nodes in a molecular graph. We include both numerical and categorical properties for drugs and materials, offering a benchmark for evaluation across diverse chemical spaces. Model performance is validated across up to nine metrics, including distribution coverage, diversity, and condition control capacity for various properties.
Datasets and Input Conditions
We have one polymer dataset [thornton2012polymer] for materials, featuring three numerical gas permeability conditions: O2Perm, CO2Perm, and N2Perm. For drug design, we create three class-balanced datasets from MoleculeNet [wu2018moleculenet]: HIV, BBBP, and BACE, each with a categorical property related to HIV virus replication inhibition, blood-brain barrier permeability, or human $\beta$-secretase 1 inhibition, respectively. We have two more numerical conditions for synthesizability from synthetic accessibility (SAS) and complexity scores (SCS) [ertl2009estimation, coley2018scscore].
Table 1: Multi-Conditional Generation of 10K Polymers: Results on the synthetic score (Synth.) and three numerical properties (gas permeability for O2, N2, CO2). MAE is calculated between the input conditions and the properties of the generated polymers using Oracles. Best results are highlighted.